### v1

@@ 6.3 Analyzing the efficacy of MSVP
Analyzing the maximum singular values. We analyze the effectiveness of the proposed MSVP algorithm, which is the core of SpecFormer. Figure 2(a) shows the maximum singular values of both ViT and our SpecFormer. It indicates that with minimal additional cost, our SpecFormer effectively limits the maximum singular values across all layers, which are smaller than those of vanilla ViTs. Hence, our proposed MSVP algorithm can effectively control the maximum singular value of the attention layers, thereby ensuring stability even under extreme perturbations.
Feature visualization. We further illustrate t-SNE feature visualizations of both vanilla ViT and our SpecFormer in Figures 2(b) and 2(c), employing the t-SNE technique as introduced by (Van der Maaten and Hinton, 2008). Adversarial samples introduce subtle perturbations aimed at diverting corrupted samples away from their correct categories. In these figures, each distinct cluster represents a class. While vanilla ViTs encounter difficulties in recovering adversarial samples that have been pushed into other categories, our modified model, SpecFormer, adeptly realigns the corrupted samples with their original true classes. This contrast serves to demonstrate the significant enhancement of ViTs’ adversarial robustness by our method.
Feature distribution distance. To further support this contrast quantitatively, we compute the $\mathcal{A}$-distance (Ben-David et al., 2006), which quantifies the similarity between two distributions, with a lower value indicating a higher degree of similarity. As shown in Figure 2(b) and 2(c), SpecFormer achieves a lower $\mathcal{A}$-distance compared to vanilla ViTs, highlighting its capability to effectively counteract the distribution shift induced by adversarial perturbations. Consequently, our approach enhances the safety and robustness of ViTs. To sum up, these analyses demonstrate that our SpecFormer can both efficiently and effectively learn robust features w.r.t the adversarial attacks.
@@
Relative ViT LipsFormer L2Former LNFormer SpecFormer
@@
7 Limitations
In the realm of adversarial robustness, a challenging yet frequently discussed issue is the relationship between clean accuracy and robust accuracy. In this study, we are no exception, as our experiments have revealed the intricate balance between clean accuracy and robust accuracy through the adjustment of hyperparameters. Investigating and modeling this relationship further using theoretical approaches will be a primary focus for our future research endeavors. Furthermore, as mentioned earlier, the strict adherence to complete Lipschitz continuity in mathematical terms can prove overly stringent and potentially limit the diversity of neural network representations, consequently affecting performance. Therefore, exploring theoretical models that address the tradeoff between Lipschitz continuity and expressiveness is also a worthy avenue for inquiry. Last but not least, it is essential to note that our method has only undergone validation against a limited number of attack scenarios. Conducting additional assessments across a broader spectrum of attacks, including but not limited to AutoAttack (Croce and Hein, 2020), backdoor attacks (Li et al., 2022), and BIM (Feinman et al., 2017), is imperative to provide a more comprehensive evaluation.
@@ 8 Conclusion
This paper presents a theoretical analysis of the self-attention mechanism in ViTs and establishes a relationship between the adversarial robustness of ViTs and Lipschitz continuity theory. We demonstrate that the Lipschitz continuity of ViTs in the vicinity of the input can be enhanced by penalizing the maximum singular values of the attention weight matrices. Extensive experiments validate the effectiveness of our approach, showcasing superior performance in terms of both clean and robust accuracy against various adversarial attacks.
Appendix A General Optimization Objective for MSVP
@@
where $\lambda_{q},\lambda_{k}$ and $\lambda_{v}$ are trade-off parameters. The original Transformer Vaswani et al. [2017] adopts multi-head self-attention to jointly attend to information from different subspaces at different positions. In line with that spirit, MSVP can also be added in a multi-head manner. Incorporating the summation over all the heads and layers, the overall training objective with multi-head MSVP is:
@@
where $\mathbf{W}_{l}^{*,h}$ denotes the $h^{th}$ head in the $l^{th}$ attention layer, and *** could be $Q,K$ or $V$. By restricting the maximum size of the perturbation’s largest singular value, we can control the range of the output changes when attacked, leading to a model with better robustness.The computation of maximum singular values can be seamlessly integrated into the forward process.
Appendix B Proof of Theorem 4.3
@@
Theorem 3.2 (Local Lipschitz Constant of self-attention layer).
@@
Proof.
To begin, we introduce some fundamental notations and re-formulate the self-attention mechanism as follows:
@@
Combining equation 28, 34, we have
@@
Lemma B.1 (Relationship between the spectral norm of a block row and the block matrix, Kim et al. [2021]).
@@
Utilizing this lemma, and consider the input around $\mathbf{X}_{0}$: $B_{2}\left(\mathbf{X}_{0},\delta_{0}\right):=\left\{\mathbf{X}:\left\|\mathbf{X}-\mathbf{X}_{0}\right\|_{F}\leq\delta_{0}\right\}$, we can focus on derive the spectral norm of the $i^{th}$ block row $\left[\mathbf{J}_{i1},\cdots,\mathbf{J}_{iN}\right]$ in the neighbourhood. Considering that all inputs received by ViT are bounded (for instance, all entries of image data fall within the range of 00 to $255$), and the inputs entering the Attention layer have been normalized by LayerNorm to a specific range $[0,1]$, we can further replace the norm $\|\mathbf{X}\|$ with a constant $B=\max_{\mathbf{X}\in\mathcal{X}}\|\mathbf{X}\|_{2}$, where $\mathcal{X}$ represents a bounded open set in the Euclidean space $\mathbb{R}^{N\times D}$.
@@
The first equality is in accordance with equation 37,38, while the second inequality arises from the Cauchy-Schwarz inequality and the boundedness of the input space $\mathcal{X}$ and the perturbation $\delta_{0}$. Furthermore, by utilizing $\left\|\mathbf{P}^{(i)}\right\|_{2}\leq 1$ and $\left\|\mathbf{E}_{ij}\right\|_{2}\leq 1$, we can omit them in the inequality, leave with pure weight matrices,
@@
By directly summing across the row index or utilizing the lemma B.1 can yield our main theorem.
@@
Appendix C Proof of Theorem 5.1
Theorem 4.1 (Convergence guarantee of the power iteration method [Mises and Pollaczek-Geiringer, 1929]).
Assuming the dominant singular value $\sigma_{max}(\mathbf{A})$ (which is also the eigenvalue of matrix $\mathbf{A}^{\top}\mathbf{A}$) is strictly greater than the subsequent singular values and that $\bm{u}_{0}$ is initially selected at random, then there is a probability of $1$ that $\bm{u}_{0}$ will have a non-zero component in the direction of the eigenvector linked with the dominant singular value. Consequently, the convergence will be geometric with a ratio of $\left|\frac{\sigma_{2}(\mathbf{A})}{\sigma_{max}(\mathbf{A})}\right|$.
Proof.
@@
On the other hand, the vector for iterative step $k$ can be written as $\bm{u}_{k}=\frac{A^{k}\bm{u}_{0}}{\left\|A^{k}\bm{u}_{0}\right\|}$. Combining these two equations above, we obtain $\bm{u}_{k}\to C\mathbf{v}_{1}$ as $k\rightarrow\infty$, where $C$ is a constant. Therefore, we can use the power iteration method to approximate the largest singular value. The convergence is geometric, with a ratio of $\left|\frac{\sigma_{2}}{\sigma_{1}}\right|$. ∎
Appendix D Comparison with existing bounds
Existing literature on introducing Lipschitz continuity into Transformer models comprises three notable articles[Kim et al., 2021, Dasoulas et al., 2021, Qi et al., 2023], we abbreviate their proposed modified Transformer as L2Former[Kim et al., 2021], LNFormer[Dasoulas et al., 2021], and LipsFormer[Qi et al., 2023]. We summarize the three bounds they propose for their modified attention mechanisms below. We have also included the vanilla self-attention mechanism, along with these three modified attention mechanisms and our optimization objectives, in Table 4, where we examine whether these studies address robustness and whether they simplify the analysis of mechanisms.
Table 4: Comprehensive Comparison of the Baseline Models.
Model Proposed Mechanism Lipschitz Continuity Robustness Simplified
Transformer [Dosovitskiy et al., 2020] $\operatorname{softmax}\left(\frac{\mathbf{X}\mathbf{W}^{Q}\left(\mathbf{X}\mathbf{W}^{K}\right)^{\top}}{\sqrt{D}}\right)\mathbf{X}\mathbf{W}^{V}$-
L2Former [Kim et al., 2021] $\exp\left(-\frac{\left\|\mathbf{x}_{i}^{\top}W^{Q}-\mathbf{x}_{j}^{\top}W^{K}\right\|_{2}^{2}}{\sqrt{D/H}}\right)$
LNFormer [Dasoulas et al., 2021] $\frac{Q^{\top}K}{\max\{uv,uw,vw\}}$
LipsFormer [Qi et al., 2023] $\mathbf{q}_{i}=\frac{\left(\mathbf{x}_{i}^{\top}\mathbf{W}^{Q}\right)^{\top}}{\sqrt{\left\|\mathbf{x}_{i}^{\top}\mathbf{W}^{Q}\right\|^{2}+\epsilon}}$
@@ Appendix D Comparison with existing bounds
In [Kim et al., 2021], it was demonstrated that the self-attention mechanism, due to the potential unboundedness of its inputs, is not globally Lipschitz continuous. Instead, they introduced an L2 self-attention mechanism based on the L2 distance $\left\|\mathbf{x}_{i}^{\top}W^{Q}-\mathbf{x}_{j}^{\top}W^{K}\right\|_{2}^{2}$. However, their proposed L2 attention mechanism also lacks global Lipschitz continuity in cases where $\mathbf{W}^{K}\neq\mathbf{W}^{Q}$, possessing global Lipschitz continuity only when $\mathbf{W}^{K}=\mathbf{W}^{Q}$. This constraint imposes significant limitations on various expressive capabilities of the model itself, as demonstrated by the experimental results presented in the original paper’s Appendix L. Confining the model to $\mathbf{W}^{K}=\mathbf{W}^{Q}$ results in a substantially higher loss during convergence compared to the unconstrained model, rendering the model suboptimal.
@@
In [Dasoulas et al., 2021], the authors proposed a normalization approach, denoted as $g(\mathbf{X})=\frac{\tilde{g}(\mathbf{X})}{c(\mathbf{X})}$, where $g(X)$ represents the score function in the softmax operation, to address the issue of potentially unbounded inputs, as mentioned earlier. However, their analysis is based on a simplified version of the attention mechanism, which significantly differs from the practical application of the self-attention mechanism. Additionally, through their choice of the function $c(\mathbf{X})$, it appears that the authors have implicitly assumed that the input is bounded. They defined $c(\mathbf{X})$ as follows: $c(\mathbf{X})=\max\{\|\mathbf{Q}\|_{F}\left\|\mathbf{K}^{\top}\right\|_{(\infty,2)},\|\mathbf{Q}\|_{F}\left\|\mathbf{V}^{\top}\right\|_{(\infty,2)},\left\|\mathbf{K}^{\top}\right\|_{(\infty,2)}\left\|\mathbf{V}^{\top}\right\|_{(\infty,2)}\}$. Here, the input matrix $\mathbf{X}$ is represented as $\mathbf{X}=(\mathbf{Q}\|\mathbf{K}\|\mathbf{V})$. It’s worth noting that this choice of $c(\mathbf{X})$ seems to assume boundedness in the input data, which may not fully address the issue of unbounded inputs in practical self-attention mechanisms.
@@
In [Qi et al., 2023], the authors introduced a series of modules to replace components in the vanilla Vision Transformer that they deemed to introduce instability during training. These modifications include using CenterNorm to replace LayerNorm, employing scaled cosine similarity attention (SCSA) as an alternative to vanilla self-attention, and utilizing weighted residual shortcuts controlled by additional learnable parameters, along with spectral initialization for convolutions and feed-forward connections, all aimed at ensuring that the Lipschitz constant for each component remains below $1$ for training stability.
@@
Theorem D.1 (Comparison with Existing Global Bounds).
We summarize the global Lipschitz bounds achieved after implementing modified attention in the three aforementioned articles.
$\operatorname{Lip}_{global}(\text{L2})\leq\frac{\sqrt{N}}{\sqrt{D/H}}\left(4\phi^{-1}(N-1)+1\right)\left(\sqrt{\sum_{h}\left\|\mathbf{W}^{Q,h}\right\|_{2}^{2}\left\|\mathbf{W}^{V,h}\right\|_{2}^{2}}\right)\left\|\mathbf{W}^{O}\right\|_{2},$ (43)
[…]
Appendix E Theoretical Explanation of ARD and PRM
Inspired by the analysis of Wang et al. [2022a] and Qi et al. [2023], we found out that the success of Mo et al. [2022] can be further explained by initialized optimization procedure and a bound.
Wang et al. [2022a] explained that the performance and stability of Vision Transformer are greatly influenced by the initialization. Therefore, using warmup cite timm or effective initializtaion can greatly boost the performance Qi et al. [2023] (used spectral initialization). They all consider scale the weights to have better stability.
Qi et al. [2023] section 4.1.4 using spectral initializtaion for convolution and feed-forward connection.
$\mathbf{W}_{si}=\frac{\mathbf{W}}{\sigma_{max}(\mathbf{W})},$ (46)
where $\mathbf{W}$ is Xavier-norm initialized and $\sigma_{max}(\mathbf{W})$ is its largest eigenvalue.
While Wang et al. [2022a] after applying standard initialization, scale the weights of feed-forward networks, the value projection and the output projection of atentino layers by $0.87(N^{4}M)^{-1\frac{1}{16}}$ and set the weights of residual connections as $0.81(N^{4}M)^{1\frac{1}{16}}$ to avoid instability. Better initialization process avoid the models from stucking in poor local optima and can lead to better convergence in the follow-up iterations.
Appendix F Experiments
F.1 Hyperparameter Analysis
@@
F.2 Implementation Details
The experiments were conducted on a system equipped with an Intel(R) Xeon(R) CPU E5-2687W v4 @ 3.00GHz and NVIDIA TITAN RTX. The abbreviation ViT-S stands for the Vision Transformer [Dosovitskiy et al., 2020] Small backbone, which comprises of $8$ attention blocks with $8$ heads and an embedding dimension of $768$. The DeiT-Ti backbone, on the other hand, refers to the DeiT [Touvron et al., 2021a] Tiny model, which comprises of $12$ attention blocks with $3$ heads and an embedding dimension of $192$. Finally, ConViT-Ti refers to the ConViT [d’Ascoli et al., 2021] Tiny model, comprising of $10$ layers with $4$ heads and an embedding size of $48$.
F.3 Evaluation on ImageNet
We assess the performance of our SpecFormer model on ImageNet-1k, a widely used large-scale dataset. We conduct a comparative analysis of SpecFormer against four other baseline models under both standard training and adversarial training paradigms. For standard training, we evaluate the model’s adversarial robustness using FGSM [Goodfellow et al., 2015] and PGD-2 [Madry et al., 2018] attacks, employing an attack radius of $2/255$. In the case of adversarially-trained models, we rigorously evaluate their robustness against CW-20 [Carlini and Wagner, 2017] as well as PGD-20 and PGD-100 [Madry et al., 2018] attacks, utilizing an attack radius of $8/255$. Following the training protocols outlined in [Mo et al., 2022], we conducted adversarial training of the models for a total of $10$ epochs. Our learning rate schedule aligns with their implementation, involving a reduction of the learning rate by a factor of $10$ at the $6^{th}$ and $8^{th}$ epochs. The experimental results on ImageNet are presented in Table 6. It is evident from the table that SpecFormer outperforms the other models under both standard and adversarial training paradigms, achieving the highest standard and robust accuracy. This highlights the effectiveness and superiority of our approach on a large-scale dataset.
Model Method Standard Training Adversarial Training
Standard FGSM PGD-2 Standard CW-20 PGD-20 PGD-100
ViT-B LipsFormer [Qi et al., 2023] 65.76 20.87 3.77 45.04 18.91 21.13 20.83
L2Former [Kim et al., 2021] 77.40 37.12 6.89 51.24 24.85 27.04 26.95
LNFormer [Dasoulas et al., 2021] 50.84 25.78 0.54 30.93 11.53 14.08 14.04
Transformer [Dosovitskiy et al., 2020] 79.11 41.45 10.79 60.81 30.92 32.58 32.35
SpecFormer (Ours) 80.04 43.51 11.59 62.30 31.87 32.82 32.56

### v2

@@ 6.3 Analyzing the efficacy of MSVP
Analyzing the maximum singular values. We analyze the effectiveness of the proposed MSVP algorithm, which is the core of SpecFormer. Figure 2(a) shows the maximum singular values of both ViT and our SpecFormer. It indicates that with minimal additional cost, our SpecFormer effectively limits the maximum singular values across all layers, which are smaller than those of vanilla ViTs. Hence, our proposed MSVP algorithm can control the maximum singular value of the attention layers, ensuring stability even under extreme perturbations.
Feature visualization. We further illustrate feature visualizations of both vanilla ViT and our SpecFormer in Figures 2(b) and 2(c), employing the t-SNE technique in [31]. Adversarial attacks cause adversarial samples (orange dots) to be misclassified (evidenced by the deviation of green and orange dots within classification clusters, Figure 2(b)). Better alignment between clean points (green) and adversarial points (orange) within each cluster indicates a more robust model. It is evident that our proposed SpecFormer’s features align better than those of vanilla ViTs. This is further confirmed by the $\mathcal{A}$-distance [5], shown at the top of the images. A smaller distance signifies greater similarity between clean and attacked features, demonstrating enhanced robustness under perturbations. This contrast highlights the significant improvement in ViTs’ adversarial robustness achieved by our method.
@@
Relative TransFormer LipsFormer L2Former LNFormer SpecFormer
@@ 7 Conclusion
This paper presents a theoretical analysis of the self-attention mechanism in ViTs through the lens of adversarial robustness and Lipschitz continuity theory. We develop the local Lipschitz constant bound based on the investigation and further introduce the Maximum Singular Value Penalization (MSVP) to enhance the robustness of ViTs. Our theoretical bounds demonstrate that the Lipschitz continuity of ViTs in the vicinity of the input can be bounded by the maximum singular values of the attention weight matrices. Empirical validation across four ViT variants on four datasets, under both standard and adversarial training and various attacks (FGSM, CW, PGD, AutoAttack), demonstrates the superiority of our proposed model. The comprehensive evaluation underscores the practical applicability and robustness of our approach, making it a valuable contribution to the fields of adversarial robustness and Vision Transformer safety.
Acknowledgements
Qi WU acknowledges the support from The CityU-JD Digits Joint Laboratory in Financial Technology and Engineering and The Hong Kong Research Grants Council [General Research Fund 11219420/9043008 ]. The work described in this paper was partially supported by the InnoHK initiative, the Government of the HKSAR, and the Laboratory for AI-Powered Financial Technologies.
Appendix 0.A General Optimization Objective for MSVP
@@
where $\lambda_{q},\lambda_{k}$ and $\lambda_{v}$ are trade-off parameters. The original Transformer [53] adopts multi-head self-attention to jointly attend to information from different subspaces at different positions. In line with that spirit, MSVP can also be added in a multi-head manner. Incorporating the summation over all the heads and layers, the overall training objective with multi-head MSVP is:
@@
where $\mathbf{W}_{l}^{*,h}$ denotes the $h^{th}$ head in the $l^{th}$ attention layer, and * could be $Q,K$ or $V$. By restricting the maximum size of the perturbation’s largest singular value, we can control the range of the output changes when attacked, leading to a model with better robustness.The computation of maximum singular values can be seamlessly integrated into the forward process.
Appendix 0.B Proof of Theorem 2
@@
Theorem 3.2 1 (Local Lipschitz Constant of self-attention layer)
@@
Proof
[…]
@@
Combining equation 25, 31, we have
@@
Lemma 1 (Relationship between the spectral norm of a block row and the block matrix, [26])
@@
Utilizing this lemma, and consider the input around $\mathbf{X}_{0}$:
$B_{2}\left(\mathbf{X}_{0},\delta_{0}\right):=\left\{\mathbf{X}:\left\|\mathbf{X}-\mathbf{X}_{0}\right\|_{F}\leq\delta_{0}\right\},$
we can focus on derive the spectral norm of the $i^{th}$ block row $\left[\mathbf{J}_{i1},\cdots,\mathbf{J}_{iN}\right]$ in the neighbourhood. Considering that all inputs received by ViT are bounded (for instance, all entries of image data fall within the range of 00 to $255$), and the inputs entering the Attention layer have been normalized by LayerNorm to a specific range $[0,1]$, we can further replace the norm $\|\mathbf{X}\|$ with a constant $B=\max_{\mathbf{X}\in\mathcal{X}}\|\mathbf{X}\|_{2}$, where $\mathcal{X}$ represents a bounded open set in the Euclidean space $\mathbb{R}^{N\times D}$.
@@
The first equality is in accordance with equation 34,35, while the second inequality arises from the Cauchy-Schwarz inequality and the boundedness of the input space $\mathcal{X}$ and the perturbation $\delta_{0}$. Furthermore, by utilizing $\left\|\mathbf{P}^{(i)}\right\|_{2}\leq 1$ and $\left\|\mathbf{E}_{ij}\right\|_{2}\leq 1$, we can omit them in the inequality, leave with pure weight matrices,
@@
By directly summing across the row index or utilizing the lemma 1 can yield our main theorem.
@@
Appendix 0.C Convergence Guarantee of Power Iteration Methods
Theorem 0.C.1 (Convergence guarantee of the power iteration method [33])
Proof
@@ Appendix 0.C Convergence Guarantee of Power Iteration Methods
On the other hand, the vector for iterative step $k$ can be written as $\bm{u}_{k}=\frac{A^{k}\bm{u}_{0}}{\left\|A^{k}\bm{u}_{0}\right\|}$. Combining these two equations above, we obtain $\bm{u}_{k}\to C\mathbf{v}_{1}$ as $k\rightarrow\infty$, where $C$ is a constant. Therefore, we can use the power iteration method to approximate the largest singular value. The convergence is geometric, with a ratio of $\left|\frac{\sigma_{2}}{\sigma_{1}}\right|$.
Appendix 0.D Comparison with existing bounds
@@
In [26], it was demonstrated that the self-attention mechanism, due to the potential unboundedness of its inputs, is not globally Lipschitz continuous. Instead, they introduced an L2 self-attention mechanism based on the L2 distance $\left\|\mathbf{x}_{i}^{\top}W^{Q}-\mathbf{x}_{j}^{\top}W^{K}\right\|_{2}^{2}$. However, their proposed L2 attention mechanism also lacks global Lipschitz continuity in cases where $\mathbf{W}^{K}\neq\mathbf{W}^{Q}$, possessing global Lipschitz continuity only when $\mathbf{W}^{K}=\mathbf{W}^{Q}$. This constraint imposes significant limitations on various expressive capabilities of the model itself, as demonstrated by the experimental results presented in the original paper’s Appendix L. Confining the model to $\mathbf{W}^{K}=\mathbf{W}^{Q}$ results in a substantially higher loss during convergence compared to the unconstrained model, rendering the model suboptimal.
@@
In [13], the authors proposed a normalization approach, denoted as $g(\mathbf{X})=\frac{\tilde{g}(\mathbf{X})}{c(\mathbf{X})}$, where $g(X)$ represents the score function in the softmax operation, to address the issue of potentially unbounded inputs, as mentioned earlier. However, their analysis is based on a simplified version of the attention mechanism, which significantly differs from the practical application of the self-attention mechanism. Additionally, through their choice of the function $c(\mathbf{X})$, it appears that the authors have implicitly assumed that the input is bounded. They defined $c(\mathbf{X})$ as follows: $c(\mathbf{X})=\max\{\|\mathbf{Q}\|_{F}\left\|\mathbf{K}^{\top}\right\|_{(\infty,2)},\|\mathbf{Q}\|_{F}\left\|\mathbf{V}^{\top}\right\|_{(\infty,2)},\left\|\mathbf{K}^{\top}\right\|_{(\infty,2)}\left\|\mathbf{V}^{\top}\right\|_{(\infty,2)}\}$. Here, the input matrix $\mathbf{X}$ is represented as $\mathbf{X}=(\mathbf{Q}\|\mathbf{K}\|\mathbf{V})$. It’s worth noting that this choice of $c(\mathbf{X})$ seems to assume boundedness in the input data, which may not fully address the issue of unbounded inputs in practical self-attention mechanisms.
@@
In [42], the authors introduced a series of modules to replace components in the vanilla Vision Transformer that they deemed to introduce instability during training. These modifications include using CenterNorm to replace LayerNorm, employing scaled cosine similarity attention (SCSA) as an alternative to vanilla self-attention, and utilizing weighted residual shortcuts controlled by additional learnable parameters, along with spectral initialization for convolutions and feed-forward connections, all aimed at ensuring that the Lipschitz constant for each component remains below $1$ for training stability.
@@
Appendix 0.E Experiments
0.E.1 Hyperparameter Analysis
@@
0.E.2 Implementation Details
The experiments were conducted on a system equipped with an Intel(R) Xeon(R) CPU E5-2687W v4 @ 3.00GHz and NVIDIA TITAN RTX. The abbreviation ViT-S stands for the Vision Transformer [17] Small backbone, which comprises of $8$ attention blocks with $8$ heads and an embedding dimension of $768$. The DeiT-Ti backbone, on the other hand, refers to the DeiT [50] Tiny model, which comprises of $12$ attention blocks with $3$ heads and an embedding dimension of $192$. Finally, ConViT-Ti refers to the ConViT [18] Tiny model, comprising of $10$ layers with $4$ heads and an embedding size of $48$.